The questions are rewrite the following expression using one or more properties of logarithims and evaluate without a calculator.
1. log 5 (25 2/5 -)=
Plese check and see if this is correct. log5 25 2 - log5 5, 2 log5 25- log5 5, 2x2-1= 4-1=3
2. log5 125 + log5 5=
I was able to figure out a portion of the problem but then got lost. Please check and help me with the rest. log5 (125x5)= log5 625 ( this is where I got lost)
3. log9 3 - log9 243=
I think I figured out a portion of this problem as well, but then got lost. Please check and help me with the rest. log9 3/243 reduced to log9 1/81,
log9 81 ( this is where I got lost)

Answer by stanbon(48535)

The questions are rewrite the following expression using one or more properties of logarithims and evaluate without a calculator.
1. log 5 (25^2/5)= Log5(25^2) - log5(5) = 2log(5^2) - 1 = 4-1 = 3
----------------------------
Note the "-" to the right of the 5 has no meaning is nothing
follows it.
------------------------------
============================================================
2. log5 (125) + log5 (5) = 2 + 1 = 3
=================================================================
I was able to figure out a portion of the problem but then got lost. Please check and help me with the rest. log5 (125x5)= log5 625 ( this is where I got lost)
--
Correct answer: log5(125*5) = log5(125) + log5(5) = 3+1 = 4
============================================================
3. log9 3 - log9 243=
Correct answer:
[log3/log9] - [log243]/log9]
Using base 3 to would get:
[1/2] - [5/2] = -2
==========================
Cheers,
Stan H.